Systems Staff Meeting Notes 8/24/98

Systems Staff Meeting Notes
August 24, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Sharon Henry, Bobbie Meng, Suzy Shaw, Leota Whitfield

  1. The Network Coordinator position was not listed in the Sunday paper as was expected.
  2. Mail macro virus - McAffee caught it in LAC. It came in a Chinese publication attachment sent to Richard Phillips via email. It is a CAP virus only associated with attachments. We have recommended to staff members having performance problems that they turn off the virus checker. We may have to make sure they have it turned on when checking their email. The virus can go across platforms.
  3. Some staff members are passwording access to their machines. If the password is unknown or lost, it would require shorting pins to regain access to the machine. There is no library wide policy on this situation. MSL machines are all passworded, but a master list is maintained. Bill will check with the Directors at the next LMG meeting. Need to make sure that shorting the pins works on Gateways. The CMOS should be passworded with our universal password. The problem is similar with screen saver passwords.
  4. John Ingram has had 2 folders disappear from the mail server over the last two weeks. Recovery is difficult at this point in time until we get the upgrade.
  5. Erich has reported incidents of machine tampering in Preservation.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Nothing major has been ordered or received within the past week.
  2. CISCO switch swapped out, waiting on router pickup.
  3. Epson printer waiting for Scott to get the RMA # and return it.
  1. Stacy Wall is experiencing a font problem with Outlook on Dale's machine.
  1. We got the quote from Digital for the fiber runs for the classroom, $3600. Funding is coming from Tigert for this - no time estimate yet. This will enable the switch move.
Software Projects Status
  1. Sharon ran the Student Update - the expiration date adds a year to the job - it ran into the year 2000. FCLA is already looking at the problem.
  2. John Humphrey's Key database has been successfully moved to NT4. Winston also did data modifications - John can no longer enter O's, only zeros.
  3. First LASSIE instruction meeting for Special Collections will be on Thursday at 9:00 AM. Brian is working on reports for Carolyn - allowing the selector to select, Resource Services to update, Administration to make changes. Carolyn wants to be able to track license expiration & IP's.
  4. USNP - Acquisitions/Serials involved in national project. Winston is working on a multi-table format.
DP Coordinator
  1. Students returning this week - hope to have schedule stabilized by Friday.
  2. Monthly Compustat tape was returned due to errors. Currently waiting on the replacement tape.
  3. Expecting yearly CRSP tapes in September.
  1. Sharon out today PM. (Aug.24)
  2. Lee out Wed. PM. (Aug.26)
  3. Bobbie out Thurs. PM. (Aug. 27)
  4. Bill & Suzy out Thurs. PM - Fri. AM. (Aug.27-28)
  5. Debra out Friday, August 28.
  6. Lee out September 1.
  7. Labor Day September 7.
  8. Sharon out September 8.
  9. Lee out Sept. 17 - on vacation, then on to retirement.

Back to the Systems Homepage

Last updated February 24, 2000
by Debra Harris